OpenStudy (bizzeizzy):

Let f(x)=x-3 and g(x)=x^2. Find f(g(4))

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

always work from the inside out \[g(4)=4^2=16\] \[f(16)=16-3\]